# Neural Incremental Input-to-State Stable Control Lyapunov Functions for Unknown Continuous-time Systems

### v1

@@ 2.3 Control Barrier Function
Definition 2.7 ([2])
Given a ct-CS $\Xi$ with compact state space ${\mathbf{X}}$. Let a $\mathcal{L}_{dh}$-smooth function, which is also Lipschitz continuous with Lipschitz constant $\mathcal{L}_{h},h:{\mathbf{X}}\rightarrow{\mathbb{R}}$ is given as
@@
Then, $h$ is said to be a control barrier function (CBF) for the system $\Xi$ in (1) if there exists a control input $u\in{\mathbf{U}}$ and class $\mathcal{K}_{\infty}$ function $\mu$ such that the following condition holds:
@@
Now, based on Definition 2.7, we aim to design the forward invariant controller $u:=g(x)$ that will make the state-space ${\mathbf{X}}$ control forward invariant. The lemma below allows us to synthesize the controller, ensuring forward invariance.
@@
Now, we assume that there exists a controller $g:{\mathbf{X}}\rightarrow{\mathbf{U}}$ such that the condition (5) is satisfied. Then, one can infer at the boundary of ${\mathbf{X}}$, the derivative of $h$ is always positive, enforcing the trajectories of the system under the controller $g$ to move towards inside the state space. Hence, the state space ${\mathbf{X}}$ becomes control forward invariant. This completes the proof. $\square$
@@ 2.4 Problem Formulation
Now, we formally define the main problem of this paper. This paper deals with the continuous-time control system, while the system map $f:{\mathbf{X}}\times{\mathbf{U}}\rightarrow{\mathbf{X}}$ being unknown. The main problem of the paper is stated below.
@@
Given a continuous-time control system $\Xi=({\mathbf{X}},{\mathbf{U}},f)$ as defined in (1) with compact state-space ${\mathbf{X}}$ and unknown dynamics $f$, the primary objective of the paper is to synthesize a forward invariant feedback controller $g:{\mathbf{X}}\times{\mathbf{W}}\rightarrow{\mathbf{U}}$ enforcing the closed-loop system $\Xi_{g}=({\mathbf{X}},{\mathbf{W}},{\mathbf{U}},f,g)$ in (2) to be incrementally input-to-state stable with respect to external input ${\mathsf{w}}\in{\mathbf{W}}$ within the state space ${\mathbf{X}}$.
@@
In contrast to previous studies on controller design [41]-[38], which depend on precise knowledge or a specific structure of the system dynamics, our objective is to develop a controller that achieves $\delta$-ISS for the closed-loop system without requiring exact knowledge or a defined structure of the dynamics.
@@ 3 Neural $\delta$-ISS Control Lyapunov Function
To solve Problem 2.9, we now try to design $\delta$-ISS-CLF, such that the closed-loop system becomes incrementally input-to-state stable according to Theorem 2.6. To do so, we first reframe the conditions of Definition 2.5 as a robust optimization problem, alongside the condition (5) of Definition 2.7 to ensure forward invariance of the controller:
$\min_{[\eta,d]}\quad$ $\eta$
s.t. $\forall x,\hat{x}\in{\mathbf{X}}:$
$\quad-V(x,\hat{x})+\alpha_{1}(|x-\hat{x}|)\leq\eta,$ (6a)
$\quad V(x,\hat{x})-\alpha_{2}(|x-\hat{x}|)\leq\eta,$ (6b)
$\forall x,\hat{x}\in{\mathbf{X}},\forall w,\hat{w}\in{\mathbf{W}}:$
$\quad\frac{\partial V}{\partial x}f(x,g(x,w))+\frac{\partial V}{\partial\hat{x}}f(\hat{x},g(\hat{x},\hat{w}))+\kappa V(x,\hat{x})-\sigma(|w-\hat{w}|)\leq\eta,$ (6c)
@@
$\quad-\frac{\partial h}{\partial x}(f(x,g(x,w)))-\mu(h(x))\leq\eta,$ (6d)
$d=[V,g,\alpha_{1},\alpha_{2},\kappa,\sigma,\mu].$
Note that $V(\cdot,\cdot)\in\{V|V:{\mathbf{X}}\times{\mathbf{X}}\rightarrow{\mathbb{R}}_{0}^{+}\},g(\cdot,\cdot)\in\{g|g:{\mathbf{X}}\times{\mathbf{W}}\rightarrow{\mathbf{U}}\},\alpha_{1},\alpha_{2},\mu\in\mathcal{K}_{\infty},\sigma\in\mathcal{K},{\kappa\in{\mathbb{R}}^{+}}$. The optimal solution of ROP $\eta^{*}\leq 0$ leads to the satisfaction of conditions of Definition 2.5 and 2.7, and thus the function $V(x,\hat{x})$ will be a valid $\delta$-ISS-CLF for the unknown system enforcing the system to be $\delta$-ISS under the forward invariant controller $g$.
However, there are several challenges in solving the ROP. They are listed as follows:
@@
    Since the structures of the controller and the $\delta$-ISS-CLF are unknown, the solution to the ROP becomes challenging.
@@
    The system dynamics $f$ is unknown, therefore, we cannot directly leverage the conditions (6c) and (6d). So, solving the ROP becomes non-trivial.
@@
    As the structures of the $\mathcal{K}_{\infty}$ functions $\alpha_{1},\alpha_{2},\sigma,\mu$ and the constant $\kappa$ are unknown, the optimization problem has become a non-convex problem.
@@
    State space is continuous in nature, so there will be infinitely many equations in the ROP, making the solution of the ROP intractable.
@@ 3.1 Architecture of neural networks
For a ct-CS $\Xi$ as in (1), the $\delta$-ISS-CLF neural network consists of an input layer with $2n$ neurons, where $n$ is the system dimension, and an output layer with one neuron, signifying the scalar output of the Lyapunov function. The network consists of $l_{v}$ hidden layers, with each hidden layer containing $h_{v}^{i},i\in[1;l_{v}]$ neurons, where both values are arbitrarily chosen.
The $\delta$-ISS-CLF is a feed-forward neural network and satisfying the condition (ii) of Definition 2.5 requires the computation of $\frac{\partial V}{\partial x}$, which necessitates a smooth activation function $\varphi(\cdot)$ (for example, Tanh, Sigmoid, Softplus etc.), facilitating derivation of smooth Jacobian values upon differentiation [23]. Hence, the resulting Neural network function is obtained by applying the activation function recursively and is denoted by:
[…]
where $\phi_{i}:{\mathbb{R}}^{h_{v}^{i}}\rightarrow{\mathbb{R}}^{h_{v}^{i}}$ is defined as $\phi_{i}(q^{i})=[\varphi(q_{1}^{i}),\ldots,\varphi(q_{h_{v}^{i}}^{i})]^{\top}$ with $q^{i}$ denoting the concatenation of outputs $q_{j}^{i},j\in[1;h_{v}^{i}]$ of the neurons in $i$-th layer. The notation for the controller $g_{\bar{\theta},\bar{b}}$ is similar and is also considered a feed-forward neural network. In this case, the input layer and the output layer have dimensions of $n+p$ and $m$, respectively. The number of hidden layers of the controller neural network is $l_{c}$ and each layer has $h_{c}^{i},i\in[1;l_{c}]$ neurons. The activation function for the controller neural network is also a slope-restricted function, but it does not need to be smooth, unlike the $\delta$-ISS-CLF neural network. The resulting controller neural network is given by,
$\begin{cases}z^{0}=[x^{\top},w^{\top}]^{\top},x\in{\mathbb{R}}^{n},w\in{\mathbb{R}}^{p}\\ z^{i+1}=\phi_{i}(\bar{\theta}^{i}z^{i}+\bar{b}^{i})\hskip 1.99997pt\text{for}\hskip 1.99997pti\in[0;l_{c}-1],\\ g_{\bar{\theta},\bar{b}}(x,w)=\bar{\theta}^{l_{c}}z^{l_{c}}+\bar{b}^{l_{c}}.\end{cases}$
@@ 3.2 Formal Verification Procedure of $\delta$-ISS-CLF
Moving onto challenge (C2) in ROP (6), we make the following assumption.
@@
We consider having access to the black box or simulator model of the system. Hence, given a state-input pair $(x,u)$, we will be able to generate the next state $f(x,u)$.
This assumption allows us to use conditions (6c) and (6d) without requiring explicit system knowledge. Now, to overcome the challenge (C3), another assumption is considered.
@@
We consider that class $\mathcal{K}_{\infty}$ functions $\alpha_{i},i\in\{1,2\}$ are of degree $\gamma_{i}$ with respect to $|x-\hat{x}|$ and class $\mathcal{K}_{\infty}$ function $\sigma$ is of degree $\gamma_{w}$ with respect to $|w-\hat{w}|$, i.e., $\alpha_{i}(|x-\hat{x}|)=k_{i}|x-\hat{x}|^{\gamma_{i}}$, and $\sigma(|w-\hat{w}|)=k_{w}|w-\hat{w}|^{\gamma_{w}}$, where $k:=[k_{1},k_{2},k_{w}],\gamma:=[\gamma_{1},\gamma_{2},\gamma_{w}]$ are user-specific constants. Additionally, the class $\mathcal{K}_{\infty}$ function $\mu$ in (6d) is considered to be of the form $\mu_{h}h(x),\mu_{h}\in{\mathbb{R}}^{+}$, while the constants $\kappa,\mu_{h}\in{\mathbb{R}}^{+}$ are also user-defined.
Now, to avoid the infinite number of equations in the ROP, we leverage a sampling-based approach to get samples from the compact state space ${\mathbf{X}}$ as well as the input space ${\mathbf{W}}$, and convert the ROP in (6) into a scenario convex program (SCP) [3]. Note that these datasets will help with the training procedure of the neural networks, which will be discussed in the next section. We collect samples $x_{s}$ and $w_{p}$ from ${\mathbf{X}}$ and ${\mathbf{W}}$, where $s\in[1;N],p\in[1;M]$. Consider ball $B_{\varepsilon_{x}}(x_{s})$ and $B_{\varepsilon_{u}}(w_{p})$ around each sample $x_{s}$ and $w_{p}$ with radius $\varepsilon_{x}$ and $\varepsilon_{u}$, such that, ${\mathbf{X}}\subseteq\bigcup_{s=1}^{N}B_{\varepsilon_{x}}(x_{s})$ and ${\mathbf{W}}\subseteq\bigcup_{p=1}^{M}B_{\varepsilon_{u}}(w_{p})$ with:
$|x-x_{s}|$ $\leq\varepsilon_{x},\forall x\in B_{\varepsilon_{x}}(x_{s}),$ (7a)
$|w-w_{p}|$ $\leq\varepsilon_{u},\forall u\in B_{\varepsilon_{u}}(w_{p}).$ (7b)
We consider $\varepsilon=\max(\varepsilon_{x},\varepsilon_{u})$. Collecting the data points obtained upon sampling the state-space ${\mathbf{X}}$ and input space ${\mathbf{W}}$, we form the training datasets denoted by:
$\mathcal{X}$ =$=\{x_{s}|x_{s}\in{\mathbf{X}},\forall s\in[1;N]\},$ (8a)
$\mathcal{W}$ =$=\{w_{p}|w_{p}\in{\mathbf{W}},\forall p\in[1;M]\}.$ (8b)
Utilizing the data sets and following the assumptions, we construct a scenario convex problem (SCP) corresponding to the RCP (6) to alleviate the challenge (C4), defined as below:
$\min_{\eta}$ $\quad\eta$
s.t. $\forall x_{q},x_{r}\in\mathcal{X}:$
$\quad-V_{\theta,b}(x_{q},x_{r})+k_{1}|x_{q}-x_{r}|^{\gamma_{1}}\leq\eta,$ (9a)
$\quad V_{\theta,b}(x_{q},x_{r})-k_{2}|x_{q}-x_{r}|^{\gamma_{2}}\leq\eta,$ (9b)
$\forall x_{q},x_{r}\in\mathcal{X},\forall w_{q},w_{r}\in\mathcal{U}:$
$\quad\frac{\partial V_{\theta,b}}{\partial x_{q}}f(x_{q},g_{\bar{\theta},\bar{b}}(x_{q},w_{q}))+\frac{\partial V_{\theta,b}}{\partial x_{r}}f(x_{r},g_{\bar{\theta},\bar{b}}(x_{r},w_{r}))+\kappa V_{\theta,b}(x_{q},x_{r})-k_{w}|w_{q}-w_{r}|^{\gamma_{w}}\leq\eta,$ (9c)
$\forall x_{q}\in\mathcal{X},w_{q}\in{\mathcal{W}}:$
$\quad-\frac{\partial h}{\partial x}\Bigr{\rvert}_{x=x_{q}}(f(x_{q},g_{\bar{\theta},\bar{b}}(x_{q},w_{q})))-\mu_{h}h(x_{q})\leq\eta.$ (9d)
Given the finite number of data samples, the SCP involves a finite set of equations, making its solution computationally tractable. Let $\eta_{S}^{*}$ denote the optimal solution of the SCP. To establish that this solution is also feasible for the proposed ROP, we introduce the following assumptions concerning Lipschitz continuity.
@@
The function $f$ in (2) is Lipschitz continuous with respect to $x$ and $u$ over the state space ${\mathbf{X}}$ and the input space ${\mathbf{W}}$ with Lipschitz constant $\mathcal{L}_{x}$ and $\mathcal{L}_{u}$.
Since the model is unknown, the Lipschitz constants of the system can be estimated following a similar procedure as in [25, Algorithm 2].
@@
Since, the sets ${\mathbf{X}}$ and ${\mathbf{W}}$ are compact, the class $\mathcal{K}_{\infty}$ functions are Lipschitz continuous with Lipschitz constants $\mathsf{L}_{1},\mathsf{L}_{2}$ and $\mathsf{L}_{u}$, respectively, with respect to $|x-\hat{x}|$ and $|w-\hat{w}|$. The values can be estimated using the values of $k$ and $\gamma$. In addition, the Lipschitz constant $\mathcal{L}_{h}$ of the function $h$ is already known as the function $h$ is predefined.
@@
Under Assumption 3, 4 and 5, the following theorem outlines the connection of the solution of SCP (9) to that of ROP (6), providing a formal guarantee to the obtained $\delta$-ISS-CLF satisfying the incremental stability conditions under the controller $g_{\bar{\theta},\bar{b}}$.
@@
Consider a continuous-time control system (2) with compact state-space ${\mathbf{X}}$ and input space ${\mathbf{W}}$. Let $V_{\theta,b}$ be the neural $\delta$-ISS control Lyapunov function and $g_{\bar{\theta},\bar{b}}$ be the corresponding controller. Consider the optimal value of the SCP (9), $\eta_{S}^{*}$, computed using data samples collected from the state-space and input-space. Under the Assumption 3, 4 and 5, if
$\eta_{S}^{*}+\mathcal{L}\varepsilon\leq 0,$ (10)
where $\varepsilon=\max(\varepsilon_{x},\varepsilon_{u})$ as defined in (7) and $\mathcal{L}=\max\{\sqrt{2}\mathcal{L}_{L}+2\mathsf{L}_{1},\sqrt{2}\mathcal{L}_{L}+2\mathsf{L}_{2},\sqrt{2}\kappa\mathcal{L}_{L}+2L_{u}+2(\mathcal{M}_{f}\mathcal{L}_{dL}+\mathcal{M}_{L}(\mathcal{L}_{x}+\sqrt{2}\mathcal{L}_{u}\mathcal{L}_{C})),\mathcal{M}_{f}\mathcal{L}_{dh}+\mathcal{M}_{h}(\mathcal{L}_{x}+\sqrt{2}\mathcal{L}_{u}\mathcal{L}_{C})+\mu_{h}\mathcal{L}_{h}\}$, it essentially follows that the $\delta$-ISS control Lyapunov function obtained by solving the SCP ensures that the closed-loop system is incrementally input-to-state stable under the action of the controller $g_{\bar{\theta},\bar{b}}$.
@@
The proof can be found in Appendix B. $\square$

### v2

@@ 2.3 Control Barrier Function
Definition 2.7 ([ames2019control])
Given a ct-CS $\Xi=({\mathbf{X}},{\mathbf{U}},f)$ with compact state and input sets $\mathbf{X}$} and $\mathbf{U}$}. Let a $\mathcal{L}_{dh}$-smooth function, which is also Lipschitz continuous with Lipschitz constant $\mathcal{L}_{h},h:{\mathbf{X}}\rightarrow{\mathbb{R}}$ is given as
@@
Then, $h$ is said to be a control barrier function (CBF) for the system $\Xi$ in (2) if for any $(x,w)\in{\mathbf{X}}\times{\mathbf{W}}$, there exists some control input $u\in{\mathbf{U}}$ such that the following condition holds:
@@
where $\mu$ is a class $\mathcal{K}_{\infty}$ function.
Now, based on Definition 2.7, we aim to design the forward invariant controller $u:=g(x,w)$ that will make the state-space $\mathbf{X}$} control forward invariant. The lemma below allows us to synthesize the controller, ensuring forward invariance.
@@
Now, we assume that there exists a controller $g:{\mathbf{X}}\rightarrow{\mathbf{U}}$ such that condition (5) is satisfied. Then, one can infer at the boundary of $\mathbf{X}$}, the derivative of $h$ is always positive, enforcing the trajectories of the system under the controller $g$ to move towards inside the state space. Hence, the state space $\mathbf{X}$} becomes control forward invariant. This completes the proof. $\square$
@@ 2.4 Problem Formulation
Now, we formally define the main problem of this paper.
@@
Given a continuous-time control system $\Xi=({\mathbf{X}},{\mathbf{U}},f)$ as defined in (1) with compact state-space $\mathbf{X}$} and unknown dynamics $f:{\mathbf{X}}\times{\mathbf{U}}\rightarrow{\mathbf{X}}$, the primary objective of the paper is to synthesize a forward invariant feedback controller $g:{\mathbf{X}}\times{\mathbf{W}}\rightarrow{\mathbf{U}}$ enforcing the closed-loop system $\Xi_{g}=({\mathbf{X}},{\mathbf{W}},{\mathbf{U}},f,g)$ in (2) to be incrementally input-to-state stable with respect to external input ${\mathsf{w}}\in{\mathbf{W}}$ within the state space $\mathbf{X}$}.
@@
In contrast to previous studies on controller design [zamani2011backstepping]-[zaker2024controller], which depend on precise knowledge or a specific structure of the system dynamics, our objective is to develop a controller that achieves $\delta$-ISS for the closed-loop system without requiring exact knowledge or a defined structure of the dynamics.
@@ 3 Neural $\delta$-ISS Control Lyapunov Function
To solve Problem 2.9, we now try to design $\delta$-ISS-CLF, such that the closed-loop system becomes incrementally input-to-state stable according to Theorem 2.6. To do so, we first provide the following lemma:
Lemma 3.1
The closed-loop system in (2) is assured to be $\delta$-ISS within the compact state-space $\mathbf{X}$} corresponding to the input-space $\mathbf{W}$} if there exists some $V$ and $g$ that satisfies the following conditions with $\eta\leq 0$:
$\forall x,\hat{x}\in{\mathbf{X}},x=\hat{x}:\ V(x,\hat{x})=0,$ (6a)
$\forall x,\hat{x}\in{\mathbf{X}},x\neq\hat{x}:\ -V(x,\hat{x})+\alpha_{1}(|x-\hat{x}|)\leq\eta,$ (6b)
$\forall x,\hat{x}\in{\mathbf{X}},x\neq\hat{x}:\ V(x,\hat{x})-\alpha_{2}(|x-\hat{x}|)\leq\eta,$ (6c)
$\forall x,\hat{x}\in{\mathbf{X}},x\neq\hat{x},\forall w,\hat{w}\in{\mathbf{W}}:$
$\quad\frac{\partial V}{\partial x}f(x,g(x,w))+\frac{\partial V}{\partial\hat{x}}f(\hat{x},g(\hat{x},\hat{w}))+\kappa V(x,\hat{x})-\sigma(|w-\hat{w}|)\leq\eta,$ (6d)
@@
$\quad-\frac{\partial h}{\partial x}(f(x,g(x,w)))-\mu(h(x))\leq\eta.$ (6e)
Proof:
With $\eta\leq 0$, the condition (6b) and (6c) ensure that the $\delta$-ISS-CLF is bounded by the class $\mathcal{K}_{\infty}$ functions. Satisfaction of the condition (6d) with $\eta\leq 0$ enforces condition (ii) of Definition 2.5 is satisfied. Finally, with $\eta\leq 0$, satisfaction of condition (6e) implies the forward invariance of the set $\mathbf{X}$} under the controller $g$. Thereby, with $\eta\leq 0$, satisfaction of the condition (6) ensures $V$ to be a valid $\delta$-ISS-CLF. $\square$
Note that $V(\cdot,\cdot)\in\{V|V:{\mathbf{X}}\times{\mathbf{X}}\rightarrow{\mathbb{R}}_{0}^{+}\},g(\cdot,\cdot)\in\{g|g:{\mathbf{X}}\times{\mathbf{W}}\rightarrow{\mathbf{U}}\},\alpha_{1},\alpha_{2},\mu\in\mathcal{K}_{\infty},\sigma\in\mathcal{K},{\kappa\in{\mathbb{R}}^{+}}$. However, certain challenges prevent us to synthesize $V$ and $g$ using this lemma. They are listed as follows:
@@
    The structures of the controller and the $\delta$-ISS-CLF are unknown.
@@
    The system dynamics $f$ is unknown; therefore, we cannot directly leverage the conditions (6d) and (6e).
@@
    The structures of the $\mathcal{K}_{\infty}$ functions $\alpha_{1},\alpha_{2},\sigma,\mu$ and the constant $\kappa$ are unknown.
@@
    There will be infinitely many constraints in the above lemma due to the continuous state space.
@@ 3.1 Architecture of neural networks
For a ct-CS $\Xi$ as in (1), the $\delta$-ISS-CLF neural network consists of an input layer with $2n$ neurons, where $n$ is the system dimension, and an output layer with one neuron, signifying the scalar output of the Lyapunov function. The network consists of $l_{v}$ hidden layers, with each hidden layer containing $h_{v}^{i},i\in[1;l_{v}]$ neurons, where both values are arbitrarily chosen. To satisfy the condition (ii) of Definition 2.5, it requires the computation of $\frac{\partial V}{\partial x}$, which necessitates a smooth activation function $\varphi(\cdot)$ (for example, Tanh, Sigmoid, Softplus, etc.). Hence, the resulting Neural network function is obtained by applying the activation function recursively and is denoted by:
$\begin{cases}t^{0}=[x^{\top},\hat{x}^{\top}]^{\top},x,\hat{x}\in{\mathbb{R}}^{n},\\ t^{i+1}=\phi_{i}(\theta^{i}t^{i}+b^{i})\hskip 1.99997pt\text{for}\hskip 1.99997pti\in[0;l_{v}-1],\\ V_{\theta,b}(x,\hat{x})=\theta^{l_{v}}t^{l_{v}}+b^{l_{v}},\end{cases}$
where $\phi_{i}:{\mathbb{R}}^{h_{v}^{i}}\rightarrow{\mathbb{R}}^{h_{v}^{i}}$ is defined as $\phi_{i}(q^{i})=[\varphi(q_{1}^{i}),\ldots,\varphi(q_{h_{v}^{i}}^{i})]^{\top}$ with $q^{i}$ denoting the concatenation of outputs $q_{j}^{i},j\in[1;h_{v}^{i}]$ of the neurons in $i$-th layer. The notation for the controller $g_{\bar{\theta},\bar{b}}$ is similar and is also considered a feed-forward neural network. In this case, the input layer and the output layer have dimensions of $n+p$ and $m$, respectively. The number of hidden layers of the controller neural network is $l_{c}$ and each layer has $h_{c}^{i},i\in[1;l_{c}]$ neurons.
@@ 3.2 Formal Verification Procedure of $\delta$-ISS-CLF
Moving onto the challenge (C2), we make the following assumption.
@@
We consider having access to the black box or simulator model of the actual system (1). Hence, given an initial state $x$ and input signal $\mathsf{u}$, we will be able to forward simulate the system trajectory at any time $t\in{\mathbb{R}}_{0}^{+}$.
This assumption allows us to use conditions (6d) and (6e) without requiring explicit system knowledge, which is described at a later stage. Now, to overcome the challenge (C3), another assumption is considered.
@@
We consider that class $\mathcal{K}_{\infty}$ functions $\alpha_{i},i\in\{1,2\}$ are of degree $\gamma_{i}$ with respect to |xx^||x-$\hat${x}| and class $\mathcal{K}_{\infty}$ function $\sigma$ is of degree $\gamma_{w}$ with respect to |ww^||w-$\hat${w}|, i.e., $\alpha_{i}(|x-\hat{x}|)=k_{i}|x-\hat{x}|^{\gamma_{i}}$, and $\sigma(|w-\hat{w}|)=k_{w}|w-\hat{w}|^{\gamma_{w}}$, where $k:=[k_{1},k_{2},k_{w}],\Gamma:=[\gamma_{1},\gamma_{2},\gamma_{w}]$ are user-specific constants, where $\Gamma$ is chosen to make these functions convex. Additionally, the class $\mathcal{K}_{\infty}$ function $\mu$ in (6e) is considered to be of the form $\mu_{h}h(x),\mu_{h}\in{\mathbb{R}}^{+}$, while the constants $\kappa,\mu_{h}\in{\mathbb{R}}^{+}$ are also user-defined.
Now, to avoid the infinite number of equations in the Lemma 3.1, we leverage a sampling-based approach to get samples from the compact state space $\mathbf{X}$} as well as the input space $\mathbf{W}$}. Note that these datasets will also help with the training procedure of the neural networks, which will be discussed in the next section. Specifically, we draw $N$ state samples $x_{s}$ from the state space $\mathbf{X}$} where $s\in[1;N]$ and define a ball $B_{\varepsilon_{x}}(x_{s})$ of radius $\varepsilon_{x}$ such that for all $x\in{\mathbf{X}}$, there exists $x_{s}$ such that $|x-x_{s}|\leq\varepsilon_{x}$. This ensures that $\bigcup_{s=1}^{N}B_{\varepsilon_{x}}(x_{s})\supset{\mathbf{X}}$. Similarly, we sample $M$ inputs $w_{p}\in{\mathbf{W}}$ with balls of radius $\varepsilon_{u}$. We consider $\varepsilon=\max(\varepsilon_{x},\varepsilon_{u})$. Collecting the data points obtained upon sampling the state-space $\mathbf{X}$} and input space $\mathbf{W}$}, we form the training datasets denoted by:
$\mathcal{X}\!=\!\{x_{s}|\!\bigcup_{s=1}^{N}\!B_{\varepsilon_{x}}(x_{s})\!\supset\!{\mathbf{X}}\},{\mathcal{W}}\!=\!\{w_{p}|\!\bigcup_{p=1}^{M}\!B_{\varepsilon_{u}}(w_{p})\!\supset\!{\mathbf{W}}\}.$ (7)
Now, we introduce the following assumptions concerning Lipschitz continuity to formulate the main theorem of this subsection.
@@
The function $f$ in (2) is Lipschitz continuous with respect to $x$ and $u$ with Lipschitz constants $\mathcal{L}_{x}$ and $\mathcal{L}_{u}$. We assume the Lipschitz constants of the dynamics are known, even if the system is unknown.
Note that these constants can also be estimated using the approach in [FV_DD, Algorithm 2].
@@
Since, the sets $\mathbf{X}$} and $\mathbf{W}$} are compact, the class $\mathcal{K}_{\infty}$ functions are Lipschitz continuous with Lipschitz constants $\mathsf{L}_{1},\mathsf{L}_{2}$ and $\mathsf{L}_{u}$, respectively, with respect to |xx^||x-$\hat${x}| and |ww^||w-$\hat${w}|. The values can be estimated using the values of $k$ and $\gamma$. In addition, the Lipschitz constant $\mathcal{L}_{h}$ of the function $h$ is already known as the function $h$ is predefined.
@@
We first define the notation $)\mathsf{L}_{f,g}V(x,\hat{x}):=\frac{\partial V_{\theta,b}}{\partial x}f(x,g_{\bar{\theta},\bar{b}}(x,w))+\frac{\partial V_{\theta,b}}{\partial\hat{x}}f(\hat{x},g_{\bar{\theta},\bar{b}}(\hat{x},\hat{w}$)). However, since we have only a black-box model (as stated in Assumption 1) to simulate the forward trajectory, we can not determine the actual value of $f(x,g_{\bar{\theta},\bar{b}}(x,w))$ for any $x\in{\mathbf{X}},w\in{\mathbf{W}}$. To tackle the issue, we approximate the quantity as
[…]
Now the proposed approximation (8) satisfies the inequality $|\widehat{\mathsf{L}}_{f,g}V(x,\hat{x})-\mathsf{L}_{f,g}V(x,\hat{x})|\leq\delta_{V}$, where $\delta_{V}\in{\mathbb{R}}^{+}$. To formally quantify the value of $\delta_{V}$, we first provide the following lemma.
Lemma 3.3
[…]
Proof:
The proof can be found in Appendix C. $\square$
[…]
@@
[…]
$|\widehat{\mathsf{L}}_{f,g}V(x,\hat{x})-\mathsf{L}_{f,g}V(x,\hat{x})|\leq\tau\mathscr{L}_{Vx}\mathcal{M}_{f},$
where $\tau$ is the sampling time.
@@
The proof can be found in Appendix F. $\square$
Now one can leverage a similar notion to approximate the quantity $\mathsf{L}_{f,g}h(x_{q}):=\frac{\partial h}{\partial x_{q}}f(x_{q},g_{\bar{\theta},\bar{b}}(x_{q},w_{q}))$ as $\widehat{\mathsf{L}}_{f,g}h(x_{q})$ satisfying the relation $|\widehat{\mathsf{L}}_{f,g}h(x_{q})-\mathsf{L}_{f,g}h(x_{q})|\leq\delta_{h}$, where the quantifier $\delta_{h}\in{\mathbb{R}}^{+}$ can be estimated via similar procedure as in Theorem 3.4. In addition, it is easy to show that $\mathsf{L}_{f,g}h(x_{q})$ is Lipschitz continuous with respect to $x$ and $u$ with Lipschitz constants $\mathscr{L}_{hx}:=\mathcal{M}_{f}\mathcal{L}_{dh}+\mathcal{M}_{h}(\mathcal{L}_{x}+\mathcal{L}_{u}\mathcal{L}_{C})$ and $\mathscr{L}_{hu}:=\mathcal{M}_{h}\mathcal{L}_{u}\mathcal{L}_{C}$ respectively. Now we state the main theorem of this section.
Theorem 3.5
The closed-loop system (2) is guaranteed to be $\delta$-ISS within the compact state-space $\mathbf{X}$} if the following conditions are satisfied by the trained networks $V_{\theta,b}$ and $g_{\bar{\theta},\bar{b}}$ with $\hat{\eta}+\mathcal{L}\varepsilon\leq 0$ where $\varepsilon=\max(\varepsilon_{x},\varepsilon_{u})$ as defined in (7) and $\mathcal{L}=\max\{\sqrt{2}\mathcal{L}_{L}+2\mathsf{L}_{1},\sqrt{2}\mathcal{L}_{L}+2\mathsf{L}_{2},\sqrt{2}\kappa\mathcal{L}_{L}+2\mathsf{L}_{u}+\mathscr{L}_{Vx}+\mathscr{L}_{Vu},\mathscr{L}_{hx}+\mathscr{L}_{hu}+\mu_{h}\mathcal{L}_{h}\}$:
$\forall x_{q},x_{r}\in\mathcal{X},x_{q}=x_{r}:V_{\theta,b}(x_{q},x_{r})=0,$ (9a)
$\forall x_{q},x_{r}\in\mathcal{X},x_{q}\neq x_{r}:-V_{\theta,b}(x_{q},x_{r})+k_{1}|x_{q}-x_{r}|^{\gamma_{1}}\leq\hat{\eta}$, (9b)
$\forall x_{q},x_{r}\in\mathcal{X},x_{q}\neq x_{r}:V_{\theta,b}(x_{q},x_{r})-k_{2}|x_{q}-x_{r}|^{\gamma_{2}}\leq\hat{\eta}$, (9c)
${\forall x_{q},x_{r}\in\mathcal{X},x_{q}\neq x_{r},\forall w_{q},w_{r}\in{\mathcal{W}}:}$
${\quad\widehat{\mathsf{L}}_{f,g}V(x_{q},x_{r})+\kappa V_{\theta,b}(x_{q},x_{r})-k_{w}|w_{q}-w_{r}|^{\gamma_{w}}+\delta_{V}\leq\hat{\eta},}$ (9d)
${\forall x_{q}\in\mathcal{X},w_{q}\in{\mathcal{W}}:-\widehat{\mathsf{L}}_{f,g}h(x_{q})-\mu_{h}h(x_{q})-\delta_{h}\leq\hat{\eta}.}$ (9e)
Proof:
The proof can be found in Appendix E. $\square$